# bibliography.bib

### v2

@@ Abstract
1 Introduction
One of the most important problems in representation theory is understanding the category of algebraic representations of the general linear group $\operatorname{GL}(n)$, as well as the limit as $n\to\infty$ of these categories – the category of polynomial functors ([macdonald_symmetric_1995],[friedlander_cohomology_1997]). While over algebraically closed fields of characteristic zero representations of $GL(n)$ are fully understood through the classical work of Schur and Weyl, in positive characteristic the situation is much more complicated, and even computing the dimensions of irreducible representations is very difficult in general. However, the classification of irreducible representations is characteristic free - they are classified by highest weights $(\lambda_{1},...,\lambda_{n})$ where $\lambda_{1}\geq...\geq\lambda_{n}$, $\lambda_{i}\in\mathbb{Z}$ ([jantzen_algebraic_groups_2003]). The story is similar for the general linear supergroups $\operatorname{GL}(m|n)$ and the corresponding categories of superpolynomial functors ([axtell_superpolynomial_2013],[drupieski_cohomology_2016]) - the category of representations of $\operatorname{GL}(m|n)$ is rather complicated already in characteristic zero and much more so in positive characteristic, but the classification of irreducible representations by highest weights is available and uniform in all characteristics ([cheng_wang_superalgebras_2012]).
@@ 1 Introduction
Finally, in Section 4, we classify the irreducible representations of general linear groups in $\operatorname{Ver}_{4}^{+}$. Namely, we first show that simple representations of $\operatorname{GL}(P)$ are indexed by integer degree, with two representations in each degree unless the degree is $2\pmod{4}$, in which there is only one simple representation. Then we prove that simple representations of $\operatorname{GL}(m\mathds{1}+nP)$ are labeled by pairs of an irreducible representation of $\operatorname{GL}(m)$ and an irreducible representation of $\operatorname{GL}(nP)$, and the latter are, in turn, labeled by $n$-tuples of irreducible $\operatorname{GL}(P)$-representations $L_{1},\dots,L_{n}$ with $\deg L_{1}\geq\cdots\geq\deg L_{n}$. We also describe certain subgroups of $\operatorname{GL}(nP)$ and classify their representations. Finally, we explicitly describe the structure and tensor products of irreducible representations of $\operatorname{GL}(P)$ and $\operatorname{GL}(\mathds{1}+P)$.
1.1 Acknowledgments
I am deeply grateful to my advisor, Pavel Etingof, for both suggesting the problems in this paper and providing a huge amount of insightful advice on how to approach both the proofs and general ways of thinking about these problems. I am also grateful to Arun Kannan for our helpful conversations about the Verlinde categories, Andrew Snowden and Karthik Ganapathy for asking many excellent questions about $\operatorname{Ver}_{4}^{+}$ during a seminar talk on this paper, some of which are answered here, and Kevin Coulembier for helpful comments.
2 Preliminaries
@@
2.1 Symmetric Tensor Categories
@@
2.2 Commutative algebras and (affine) group schemes in tensor categories
@@
$\sigma_{G}$ $=i_{G}^{*}:\mathcal{O}(G)\to\mathcal{O}(G)$
@@
$\operatorname{AffGrp}\mathcal{C}\cong(\operatorname{CommHopfAlg}\mathcal{C})^{\text{op}}$
, where $\operatorname{AffGrp}\mathcal{C}$ is the category of affine group schemes in $\mathcal{C}$.
@@
An affine group scheme is finite if $\mathcal{O}(G)$ is an object, not just an ind-object.
@@
2.2.1 The fundamental group of a symmetric tensor category
@@ 2.2.2 Distribution Algebras
2.3 The category $\operatorname{Ver}_{4}^{+}$
@@
In $\operatorname{Ver}_{4}^{+}(k)$, a commutative algebra $A$ corresponds to an ordinary, possibly non-commutative algebra over $k$ equipped with a derivation $d$ with $d^{2}=0$. Since $d$ is primitive,
@@
2.3.1 $\operatorname{Ver}_{4}^{+}$ as reduction of $\operatorname{sVec}$ to characteristic 2
In this section, let $k:=\mathbb{Q}_{2}(\sqrt{2})$, where $\mathbb{Q}_{2}$ is the field of $2$-adic numbers. We recall how $\operatorname{Ver}_{4}^{+}(\mathbb{F}_{2})$ can be constructed as a non-semisimple reduction of $\operatorname{sVec}(k)$ over $k$. This construction was described in [venkatesh_hilbert_2016], Section 2.3. Let $k\mathbb{Z}/2\mathbb{Z}=\langle 1,g\rangle$, $g^{2}=1$. Recall that in characteristic not 2, the category of supervector spaces is $\operatorname{Rep}(k[\mathbb{Z}/2\mathbb{Z}])$ with associated $R$-matrix
@@ 2.3.1 $\operatorname{Ver}_{4}^{+}$ as reduction of $\operatorname{sVec}$ to characteristic 2
2.3.2 The fundamental group of $\operatorname{Ver}_{4}^{+}$
@@
$k\pi=k[d]/d^{2}$
, the commutative and cocommutative Hopf algebra in $\operatorname{Ver}_{4}^{+}$ actually contained in $\operatorname{Vec}$. As a functor of points, it sends $A$ to elements that square to 0 in $\ker d\subset A$ under addition.
@@
It then suffices to determine $\eta_{P}\in\operatorname{GL}(P\otimes A)$ and ensure $\eta_{P\otimes P}$ respects isomorphisms $P\otimes P\cong 2P$. Writing $\eta_{P}$ as a matrix $\begin{pmatrix}a&a^{\prime}\\ b&a+b^{\prime}\end{pmatrix}$ with entries in $A$, we know that $\eta_{P}$ must be trivial on $\operatorname{im}d$, so $a^{\prime}=0,a+b^{\prime}=1$. Additionally, since $\eta$ is additive, $\eta_{nP}$ fixes $n\mathds{1}\subset nP$. For $x\in nP$ not in $\ker d$, $\eta(x\otimes 1)=x\otimes(1+b^{\prime})+x^{\prime}\otimes b$.
@@
and we can check this requires $b^{2}=0$. Then $\eta_{P\otimes P}(d\otimes 1)=(d\otimes 1)\otimes(1+b^{\prime})+(d\otimes d)\otimes b$, and likewise for $1\otimes d$. Therefore, $\eta_{P}=\begin{pmatrix}1&0\\ b&1\end{pmatrix}$ where $b^{2}=0$, $b^{\prime}=0$, so $\eta$ are parametrized by elements in $\ker d$ that square to 00.
@@
2.4 Highest Weights
Let $G$ be an affine group scheme of finite type in a symmetric tensor category $\mathcal{C}$ with $\phi:\pi_{1}(\mathcal{C})\to G$. Let $f:\mathbb{G}_{m}\to G$ be a homomorphism which lands in the centralizer of $\phi(\pi_{1}(\mathcal{C}))$. Let $Z(f)$ be the centralizer of the image of $f$ in $G$. Then for every irreducible representation $L$ of $G$ compatible with $\phi$, we have the representation $\operatorname{hw}_{f}(L)$ of $Z(f)$ compatible with $\phi$ on the eigenobject of $\mathbb{G}_{m}$ on $L$ (acting via $f$) with largest eigenvalue (which is an integer). We call $\operatorname{hw}_{f}(L)$ the highest weight of $L$ with respect to $f$.
@@ 2.4 Highest Weights
$\operatorname{Dist}(G)\cong\operatorname{Dist}(N_{-})\otimes\operatorname{Dist}(Z)\otimes\operatorname{Dist}(N_{+})$ (1)
@@
Set $B_{+}=Z\times N_{+}$, then the multiplication map is an isomorphism of ind-objects
@@
3 Examples
@@ 3.1 Symmetric algebras
$\operatorname{Sym}^{n}P\equiv\begin{cases}P&n\text{ odd}\\ \mathds{1}\oplus\mathds{1}&n\text{ even}.\end{cases}$
3.2 $\mathbb{G}_{a},\mathbb{G}_{m}$, and friends
@@
Proposition 3.5.
There are three Lie algebra structures on $L=\langle a,\alpha\rangle$, $a$ even, $\alpha$ odd, up to isomorphism:
  • $L$ abelian
  • $[a,\alpha]=\alpha$, all other brackets zero
  • $[\alpha,\alpha]=a$, all other brackets zero
@@
In particular, the Lie algebra structure on $k^{1|1}$ with $[x,\xi]=\xi$ reduces to the Lie algebra of $M_{1}$: consider $v=x+\xi$ and $w=\sqrt{2}\xi$. Then $dv=w$ and $dw=\sqrt{2}w$. Moreover, $[v,v]=0$, $[v,w]=w$, and $[w,w]=0$. Reducing modulo $\sqrt{2}$, we have $dv=w$, $dw=0$, and $[v,w]=w$ is the only nonzero bracket, which is $\mathfrak{aff}(0,1)$, the Lie algebra of the group $\operatorname{Aff}(0,1)$ of affine transformations of $k^{(0,1)}$.
@@
3.3 General linear groups and subgroups
@@
$\begin{pmatrix}m\mathds{1}\otimes(m\mathds{1})^{*}&m\mathds{1}\otimes(nP)^{*}\\ nP\otimes(m\mathds{1})^{*}&nP\otimes(nP)^{*}\end{pmatrix}$
@@
$\begin{pmatrix}F&C&C^{\prime}\\ 0&D&D^{\prime}\\ 0&E&D+E^{\prime}\end{pmatrix}$ (2)
@@
$\begin{pmatrix}D&D^{\prime}\\ E&D+E^{\prime}\end{pmatrix}=\begin{pmatrix}I_{n}&0\\ ED^{-1}&I_{n}+(ED^{-1})^{\prime}\end{pmatrix}\begin{pmatrix}D&D^{\prime}\\ 0&D\end{pmatrix}$
@@
4 Representations of general linear groups in $\operatorname{Ver}_{4}^{+}$
@@
4.1 Representations of $\mathbb{G}^{\prime}_{a}$ and $\mathbb{G}^{\prime}_{m}$

### v3

@@ Abstract.
1. Introduction
One of the most important problems in representation theory is understanding the category of polynomial representations of the general linear group $\operatorname{GL}(n)$, as well as the limit as $n\to\infty$ of these categories – the category of polynomial functors ([macdonald_symmetric_1995],[friedlander_cohomology_1997]). While over algebraically closed fields of characteristic zero representations of $GL(n)$ are fully understood through the classical work of Schur and Weyl, in positive characteristic the situation is much more complicated, and even computing the dimensions of irreducible representations is very difficult in general. However, the classification of irreducible representations is characteristic free - they are classified by highest weights $(\lambda_{1},...,\lambda_{n})$ where $\lambda_{1}\geq...\geq\lambda_{n}$, $\lambda_{i}\in\mathbb{Z}$ ([jantzen_algebraic_groups_2003]). The story is similar for the general linear supergroups $\operatorname{GL}(m|n)$ and the corresponding categories of superpolynomial functors ([axtell_superpolynomial_2013],[drupieski_cohomology_2016]) - the category of representations of $\operatorname{GL}(m|n)$ is rather complicated already in characteristic zero and much more so in positive characteristic, but the classification of irreducible representations by highest weights is available and uniform in all characteristics ([cheng_wang_superalgebras_2012]).
@@ 1. Introduction
Although the category $\operatorname{Ver}_{4}^{+}$ is incompressible and developing Lie theory in it is therefore a new problem, it does have a non-symmetric fiber functor to Vec, which allows one to do explicit computations in $\operatorname{Ver}_{4}^{+}$ involving vectors, as long as one remembers to apply the $R$-matrix to compensate for the failure of this functor to be symmetric. This approach will be heavily used in the paper.
@@
Finally, in Section 4, we classify the irreducible representations of general linear groups in $\operatorname{Ver}_{4}^{+}$. Namely, we first show that simple representations of $\operatorname{GL}(P)$ are indexed by integer degree, with two representations in each degree unless the degree is $2\pmod{4}$, in which there is only one simple representation. Then we prove that simple representations of $\operatorname{GL}(m\mathds{1}+nP)$ are labeled by pairs of an irreducible representation of $\operatorname{GL}(m)$ and an irreducible representation of $\operatorname{GL}(nP)$, and the latter are, in turn, labeled by $n$-tuples of irreducible $\operatorname{GL}(P)$-representations $L_{1},\dots,L_{n}$ with $\deg L_{1}\geq\cdots\geq\deg L_{n}$. We also describe certain subgroups of $\operatorname{GL}(nP)$ and classify their representations. We explicitly describe the structure and tensor products of irreducible representations of $\operatorname{GL}(P)$ and $\operatorname{GL}(\mathds{1}+P)$. Finally, we conjecture a Steinberg tensor product theorem for $\operatorname{GL}(m+nP)$.
1.1. Acknowledgments
I am deeply grateful to my advisor, Pavel Etingof, for both suggesting the problems in this paper and providing a huge amount of insightful advice on how to approach both the proofs and general ways of thinking about these problems. I am also grateful to Arun Kannan for our helpful conversations about the Verlinde categories and the Steinberg tensor product theorem, Andrew Snowden and Karthik Ganapathy for asking many excellent questions about $\operatorname{Ver}_{4}^{+}$ during a seminar talk on this paper, some of which are answered here, and Kevin Coulembier for our discussions on $\operatorname{Ver}_{4}^{+}$ and the Steinberg tensor product theorem. I am also grateful to the anonymous reviewers who provided feedback on this article. This work was partially supported by NSF grant DMS-2001318.
2. Preliminaries
@@
2.1. Symmetric Tensor Categories
@@
2.2. Commutative algebras and (affine) group schemes in tensor categories
@@
$\sigma_{G}$ $=i_{G}^{*}:\mathcal{O}(G)\to\mathcal{O}(G).$
@@
$\operatorname{AffGrp}\mathcal{C}\cong(\operatorname{CommHopfAlg}\mathcal{C})^{\text{op}},$
where $\operatorname{AffGrp}\mathcal{C}$ is the category of affine group schemes in $\mathcal{C}$.
@@
An affine group scheme is finite if $\mathcal{O}(G)$ is an object, not just an ind-object. An affine group scheme is of finite type if $\mathcal{O}(G)$ is finitely generated.
@@
2.2.1. The fundamental group of a symmetric tensor category
@@ 2.2.2. Distribution Algebras
Definition 2.11.
Let $H\subset G$ be an inclusion of affine group schemes. Then for an $H$-module $V$, the induced $G$-module is
$\operatorname{ind}_{H}^{G}V=\operatorname{Dist}(G)\otimes_{\operatorname{Dist}(H)}V.$
2.3. The category $\operatorname{Ver}_{4}^{+}$
@@
Notice that $\operatorname{Ver}_{4}^{+}(k)$ admits a forgetful functor to $\text{Vec}_{k}$, although this functor does not respect the braiding. Nevertheless, this means we can treat objects in $\operatorname{Ver}_{4}^{+}$ concretely as vector spaces with additional structure (an automorphism $d$ such that $d^{2}=0$), and any relations involving the braiding can be written in terms of elements of these vector spaces and the $d$-action using the $R$-matrix. Hence, we can speak of elements and bases of objects in $\operatorname{Ver}_{4}^{+}(k)$.
For example, in $\operatorname{Ver}_{4}^{+}(k)$, a commutative algebra $A$ corresponds to an ordinary, possibly non-commutative algebra over $k$ equipped with a derivation $d$ with $d^{2}=0$. Since $d$ is primitive,
@@
2.3.1. $\operatorname{Ver}_{4}^{+}$ as reduction of $\operatorname{sVec}$ to characteristic 2
In this section, let $K:=\mathbb{Q}_{2}(\sqrt{2})$, where $\mathbb{Q}_{2}$ is the field of $2$-adic numbers. We recall how $\operatorname{Ver}_{4}^{+}(\mathbb{F}_{2})$ can be constructed as a non-semisimple reduction of $\operatorname{sVec}(K)$ over $K$. This construction was described in [venkatesh_hilbert_2016], Section 2.3. Let $H:=K[\mathbb{Z}/2\mathbb{Z}]$ be the group algebra of $\mathbb{Z}/2\mathbb{Z}$ as a Hopf algebra, i.e. it is generated by $1,g$ such that $g^{2}=1$ and $\Delta(g)=g\otimes g$. Recall that in characteristic not 2, the category of supervector spaces is the $H$-modules $\operatorname{Rep}(H)$ with associated $R$-matrix
@@ 2.3.1. $\operatorname{Ver}_{4}^{+}$ as reduction of $\operatorname{sVec}$ to characteristic 2
2.3.2. The fundamental group of $\operatorname{Ver}_{4}^{+}$
@@
$k\pi=k[d]/d^{2},$
the commutative and cocommutative Hopf algebra in $\operatorname{Ver}_{4}^{+}$ which is contained in $\operatorname{Vec}$. As a functor of points, it sends $A$ to elements that square to 0 in $\ker d\subset A$ under addition.
@@
It then suffices to determine $\eta_{P}\in\operatorname{GL}(P)(A)$ and ensure $\eta_{P\otimes P}$ respects isomorphisms $P\otimes P\cong 2P$. Writing $\eta_{P}$ as a matrix $\begin{pmatrix}a&a^{\prime}\\ b&a+b^{\prime}\end{pmatrix}$ with entries in $A$, we know that $\eta_{P}$ must be trivial on $\operatorname{im}d$, so $a^{\prime}=0,a+b^{\prime}=1$. Additionally, since $\eta$ is additive, $\eta_{nP}$ fixes $n\mathds{1}\subset nP$. For $x\in nP$ not in $\ker d$, $\eta(x\otimes 1)=x\otimes(1+b^{\prime})+x^{\prime}\otimes b$.
@@
and we can check this requires $b^{2}=0$. Then
$\eta_{P\otimes P}(d\otimes 1)=(d\otimes 1)\otimes(1+b^{\prime})+(d\otimes d)\otimes b,$
and likewise for $1\otimes d$. Therefore, $\eta_{P}=\begin{pmatrix}1&0\\ b&1\end{pmatrix}$ where $b^{2}=0$, $b^{\prime}=0$, so $\eta$ is parametrized by elements in $\ker d$ that square to 00.
@@
2.4. Highest Weights
From this section onwards, we assume our symmetric tensor category $\mathcal{C}$ satisfies hypothesis 1.3.1 in [coulembier_algebraic_2023], which allows us to develop scheme theory in $\mathcal{C}$. In particular, $\operatorname{Ver}_{4}^{+}$ satisfies this hypothesis, see [coulembier_algebraic_2023] Section 8.
Let $G$ be an affine group scheme of finite type in such a symmetric tensor category $\mathcal{C}$ with $\phi:\pi_{1}(\mathcal{C})\to G$. Let $f:\mathbb{G}_{m}\to G$ be a homomorphism which lands in the centralizer of $\phi(\pi_{1}(\mathcal{C}))$. Let $Z(f)$ be the centralizer of the image of $f$ in $G$. Then for every irreducible representation $L$ of $G$ compatible with $\phi$, we have the representation $\operatorname{hw}_{f}(L)$ of $Z(f)$ compatible with $\phi$ on the eigenobject of $\mathbb{G}_{m}$ on $L$ (acting via $f$) with largest eigenvalue (which is an integer). We call $\operatorname{hw}_{f}(L)$ the highest weight of $L$ with respect to $f$.
@@ 2.4. Highest Weights
(1) $\operatorname{Dist}(G)\cong\operatorname{Dist}(N_{-})\otimes\operatorname{Dist}(Z)\otimes\operatorname{Dist}(N_{+})$
@@
Set $B_{+}=Z\times N_{+}$; then the multiplication map is an isomorphism of ind-objects
@@
Corollary 2.16.
Because irreducible $B_{+}$-modules are in bijection with irreducible $Z$-modules, we can also interpret $L\mapsto\operatorname{hw}_{f}(L)$ as an injective map $\operatorname{hw}_{f}:\operatorname{Irrep}(G,\phi)\to\operatorname{Irrep}(B_{+},\phi)$.
Proposition 2.17.
An irreducible $B_{+}$-module $V$ lies in the image of $\operatorname{hw}_{f}$ iff $\operatorname{Hom}_{B_{+}}(V,\mathcal{O}(G))$ is nonzero.
Proof.
Suppose there exists some nonzero $h\in\operatorname{Hom}_{B_{+}}(V,\mathcal{O}(G))$. Let $M$ be the $G$-module generated by $h$ in $\mathcal{O}(G)$. It will have highest weight $V$, since the $G$-action can only lower the $\mathbb{G}_{m}$-eigenvalue by the same argument as in the proof of Theorem 2.15. Because $M$ is a subobject of $\mathcal{O}(G)$, which is an ind-object, $M$ will be an honest object in $\mathcal{C}$. Therefore, $V$ lies in the image of $\operatorname{hw}_{f}$.
Conversely, suppose that $V$ lies in the image of $\operatorname{hw}_{f}$, so there exists some irreducible $G$-module $M$ with highest weight $V$. Then the matrix coefficients of $M_{V}\subset M$, the subobject of $M$ with weight $V$, will provide the $B_{+}$-invariant map $V\to\mathcal{O}(G)$. ∎
3. Examples
@@ 3.1. Symmetric algebras
$\operatorname{Sym}^{n}P\cong\begin{cases}P&n\text{ odd}\\ \mathds{1}\oplus\mathds{1}&n\text{ even}.\end{cases}$
3.2. $\mathbb{G}_{a},\mathbb{G}_{m}$, and friends
@@
The Lie algebra structure on $k^{1|1}$ with $x$ even and $\xi$ odd and nonzero bracket $[x,\xi]=\xi$ reduces to the Lie algebra of $M_{1}$. Consider $v=x+\xi$ and $w=\sqrt{2}\xi$. Then $dv=w$ and $dw=\sqrt{2}w$. Moreover, $[v,v]=0$, $[v,w]=w$, and $[w,w]=0$. Reducing modulo $\sqrt{2}$, we have $dv=w$, $dw=0$, and $[v,w]=w$ is the only nonzero bracket, which is $\mathfrak{aff}(0,1)$, the Lie algebra of the group $\operatorname{Aff}(0,1)$ of affine transformations of $k^{(0,1)}$.
@@
3.3. General linear groups and subgroups
@@
$\begin{pmatrix}m\mathds{1}\otimes(m\mathds{1})^{*}&m\mathds{1}\otimes(nP)^{*}\\ nP\otimes(m\mathds{1})^{*}&nP\otimes(nP)^{*}\end{pmatrix}.$
@@
(2) $\begin{pmatrix}F&C&C^{\prime}\\ 0&D&D^{\prime}\\ 0&E&D+E^{\prime}\end{pmatrix}$
@@
$\begin{pmatrix}D&D^{\prime}\\ E&D+E^{\prime}\end{pmatrix}=\begin{pmatrix}I_{n}&0\\ ED^{-1}&I_{n}+(ED^{-1})^{\prime}\end{pmatrix}\begin{pmatrix}D&D^{\prime}\\ 0&D\end{pmatrix}.$
@@
4. Representations of general linear groups in $\operatorname{Ver}_{4}^{+}$
@@
4.1. Representations of $\mathbb{G}^{\prime}_{a}$ and $\mathbb{G}^{\prime}_{m}$